OpenStudy (anonymous):

PLEASEEEEEEE HELPPP What is the vertex in the quadratic function y = - 4x^2 + 8x - 3 ? Write your answer as an ordered pair.

OpenStudy (anonymous):

First do x=-b/2a

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{8}{2\times -4}=1\)

OpenStudy (anonymous):

You can plug in the 1 for x and solve for y

OpenStudy (anonymous):

I GOT 1.5 ?

OpenStudy (anonymous):

not likely

OpenStudy (anonymous):

\[ y = - 4x^2 + 8x - 3\] \[-4+8-3=?\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

recheck your math

OpenStudy (anonymous):

is it -1 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You're close though

OpenStudy (anonymous):

y= -4(1)^2 - 8(1) -3

OpenStudy (anonymous):

Remember PEMDAS

OpenStudy (anonymous):

11?